Classify the equation zxx + xzyy = 0 for x > 0 and reduce it to canonical form.
1
Expert's answer
2020-10-12T18:59:13-0400

 

"z_{xx}+xz_{yy}=0"

Comparing the given equation with the equation


"Az_{xx}+Bz_{xy}+Cz_{yy}+Dz_x+Ez_y+Fz=G,"

we have

"A=1, B=0, C=x, D=0, E=0, F=0, G=0"

The discriminant is


"B^2-4AC=0-4(1)(x)=-4x<0, x>0"

The given equation is elliptic for all "x>0."


 The characteristic equations are given by 


"\dfrac{dy}{dx}=\dfrac{B-\sqrt{B^2-4AC}}{2A}=\dfrac{-\sqrt{-4x}}{2}=-i\sqrt{x}"

"\dfrac{dy}{dx}=\dfrac{B+\sqrt{B^2-4AC}}{2A}=\dfrac{\sqrt{-4x}}{2}=i\sqrt{x}"

Integrating these two equations, we get 

"iy+\dfrac{2}{3}x\sqrt{x}=c_1" and "-iy+\dfrac{2}{3}x\sqrt{x}=c_2"

Assume that


"\xi=iy+\dfrac{2}{3}x\sqrt{x}, \eta=-iy+\dfrac{2}{3}x\sqrt{x}"

Introducing the transformations "\alpha=\dfrac{1}{2}(\xi+\eta)" and "\beta=\dfrac{1}{2i}(\xi-\eta)," we obtain


"\alpha=\dfrac{1}{2}(iy+\dfrac{2}{3}x\sqrt{x}-iy+\dfrac{2}{3}x\sqrt{x})=\dfrac{2}{3}x\sqrt{x}"

"\beta=\dfrac{1}{2i}(iy+\dfrac{2}{3}x\sqrt{x}+iy-\dfrac{2}{3}x\sqrt{x})=y"

The transformed equation is obtained by first computing


"\bar{A}=A\alpha_{x}^2+B\alpha_{x}\alpha_{y}+C\alpha_{y}^2=x"


"\bar{B}=3A\alpha_{x}\beta_{x}+B(\alpha_{x}\beta_{y}+\alpha_{y}\beta_{x})+2C\alpha_{y}\beta_{y}=0"

"\bar{C}=A\beta_{x}^2+B\beta_{x}\beta_{y}+C\beta_{y}^2=x"

"\bar{D}=A\alpha_{xx}+B\alpha_{xy}+C\alpha_{yy}+D\alpha_{x}+E\alpha_{y}=\dfrac{1}{2\sqrt{x}}"

"\bar{E}=A\beta_{xx}+B\beta_{xy}+C\beta_{yy}+D\beta_{x}+E\beta_{y}=0"

"\bar{F}=F=0\ and\ \bar{G}=G=0"

Substitute in the following equation


"\bar{A}z_{\alpha\alpha}+\bar{B}z_{\alpha\beta}+\bar{C}z_{\beta\beta}+\bar{D}z_\alpha+\bar{E}z_\beta+\bar{F}z=\bar{G}"

Thus, canonical form of the given partial differential equation is


"xz_{\alpha\alpha}+xz_{\beta\beta}+\dfrac{1}{2\sqrt{x}}z_\alpha=0"

Or


"z_{\alpha\alpha}+z_{\beta\beta}=-\dfrac{1}{2x\sqrt{x}}z_\alpha"

"z_{\alpha\alpha}+z_{\beta\beta}=-\dfrac{1}{3\alpha}z_\alpha"